OpenStudy (cenaida):

Fill in the missing term: (x - 8)2 = x2 + ____ + 64. A. 4x B. 8x C. -16x D. 32x

ganeshie8 (ganeshie8):

use this :- \(\large (a-b^2) = a^2 - 2ab + b^2\)

OpenStudy (thejax):

-16x

OpenStudy (thejax):

as, @ganeshie8 said

ganeshie8 (ganeshie8):

\(\large (a-b)^2 = a^2 - 2ab + b^2\)